OpenStudy (anonymous):

write each improper fraction as a mixed number or each mixed number as an improper fraction. #1. 3/2 #4. 7/6 #6. 21/2 #15. 22/3 #16. 5/4 Thank you so much to those who help me :D

OpenStudy (anonymous):

@Luigi0210

OpenStudy (ranga):

3/2. Divide 3 by 2. You get quotient 1 and remainder 1. 3/2 = 1 and 1/2

OpenStudy (ranga):

\[\frac{ 3 }{ 2 } = 1\frac{ 1 }{ 2 }\]

OpenStudy (ranga):

\[\frac{ 7 }{ 6 } = 1\frac{ 1 }{ 6 }\]